1.    The monthly market demand curve for calculators among engineering students is given by

P = 100 – Q. where P is the price per calculator in dollars and Q is the number of calculators purchased per month. If the price is UD$ 30, how much revenue will calculator makers get each month? Find the price elasticity of demand for calculators at this point. What should calculator makers do to increase revenue?


1
Expert's answer
2021-06-21T14:33:15-0400

P=100-Q

If P= 30

Q=100-30=70

Revenue= P"\times Q"

"70\times30=2100"

PED="\frac{P}{Q} ." "\frac{∆Q}{∆P}"

No change in price and quantity

"\frac{30}{70} . 1= 0.4286"

PED = 0.4286


At this point it is inelastic


An increase in the price of the calculators will decrease the quantity of calculators demanded therefore increase in the revenue collected..i.e the price increases to $40 the demand drops to $60 calculators


Revenue collected

"60\times40=2400"

Hence revenue increased